# Follow the Energy, Find the Path: Riemannian Metrics from Energy-Based Models

### v1

@@ 4.3 Rotated Characters
At first glance, it may seem counterintuitive that trajectories following the geodesics move faster in high-density regions. However, this is consistent with minimizing the kinetic energy $\mathcal{E}$ in Eq.2, which enforces constant Riemannian speed (i.e., the quantity $||\dot{\mathbf{\gamma}}(t)||_{\mathbf{\gamma}(t)}$ is preserved along the trajectory) but not a constant Euclidean speed (i.e., $||\dot{\mathbf{\gamma}}(t)||$ is not constant). Since EBM-derived metrics assign lower Riemannian cost in high-density regions, maintaining constant Riemannian speed requires moving faster in Euclidean terms through these regions. The faster rotation near 0°, observed in Fig.5 and Fig. 3, thus reflects the lower Riemannian cost of traveling through high-density regions. These results confirm and extend our previous findings: metrics based on energy (i.e., proportional to $-\log p$) more effectively capture the curvature of the data manifold.
@@ 4.4 Animal Faces
[…]
@@
[…]
@@
Table 1: FID along geodesics for different Riemannian metrics. FID is computed at each trajectory point to assess on-manifold alignment. See Supp. F.4 for the $2$-$\sigma$ error.
@@ 5 Conclusion
We focus on conformal metrics, which scale the identity by a scalar field to encode density. However, EBM gradients (scores) could enable more expressive, anisotropic metrics [25]. Future work may explore these extensions with regularization or structural priors to ensure smoothness and scalability (See Supp.G for a discussion of limitations and Supp.H for broader impact).
Although this article is primarily methodological, it points to promising applications. One example is the mental rotation task, in which humans mentally rotate objects to match a target [62]. In such experiments, reaction times tend to decrease with training [63], suggesting that repeated exposure sharpens internal representations around training examples. These refined representations may concentrate in high-density regions, where mental transformations occur more quickly. As shown in Fig.3 and 5, our geodesics naturally accelerate in such high-density regions, echoing these psychophysical findings. Modeling mental simulation as geodesics on Riemannian manifolds shaped by a generative model offers a principled computational framework to understand human cognition. It provides a way to formalize and test the hypothesis that the human cognition relies on generative models to support flexible inference [64, 65, 66, 67, 68].
@@ B.1 Derivation of the Gradient of the EBM Log-Likelihood
[…]
@@
$-\log p_{\theta}(\mathbf{x})=E_{\theta}(\mathbf{x})+\log Z(\theta).$
@@
[…]
@@ B.2 EBM training algorithm
Input: Training dataset :$\mathcal{D}$, learning rate $\eta$, Replay Buffer $\mathcal{B}$, Langevin step size $\alpha$, noise scale $\sigma$, number of Langevin steps $L$
@@
[…]
      $\nabla_{\theta}\mathcal{L}_{\text{ML}}\approx\mathbb{E}_{\textnormal{Batch}}\left[\nabla_{\theta}E_{\theta}(\mathbf{x}^{+}_{i})-\nabla_{\theta}E_{\theta}(\mathbf{x}^{-}_{i})\right]$# Compute the ML loss
       $\mathcal{L}_{REG}(\theta)=\mathbb{E}_{\textnormal{Batch}}\left[\nabla_{\theta}E_{\theta}(\mathbf{x}^{+}_{i})^{2}+\nabla_{\theta}E_{\theta}(\mathbf{x}^{-}_{i})^{2}\right]$# Compute Regularization loss
       $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{ML}}-\eta\nabla_{\theta}\mathcal{L}_{REG}$# update parameters with gradient descent
@@ C.2 LAND metric
${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G}_{\text{LAND}}(\mathbf{x})}=(\alpha\,\textnormal{diag}(\mathbf{h}(\mathbf{x}))+\beta\mathbf{I})^{-1}\,\,\textnormal{s.t}\,\,h^{(j)}(\mathbf{x})=\sum_{i=1}^{N}(x_{i}^{(j)}\!-\!x^{(j)})^{2}\exp{\bigg{(}\!\!\!-\frac{||\mathbf{x}\!-\!\mathbf{x_{i}}||^{2}}{2\sigma^{2}}\bigg{)}}$ (12)
@@ C.3 RBF metric
$\lambda_{k}=\frac{1}{2}\bigg{(}\frac{\kappa}{2K}\sum_{k=1}^{K}||\mathbf{x}-\mathbf{\hat{x}_{k}}||\bigg{)}^{-2}$ (13)
@@ D.2 Neural networks architectures and Hyperparameters on the Circular Mixture of Gaussian Dataset
Refer to captionRefer to captiona)$E_{\theta}$ on the UCG datasetb)$E_{\theta}$ on the WCG datasetc)$-\log p_{\mathcal{M}}$ on the UCG datasetd)$-\log p_{\mathcal{M}}$ on the WCG dataset
Figure 10: Energy Landscape on the UCG and WCG datasets. (a, b) shows the energy landscape learned by the EBMs on the UCG and WCG datasets, respectively. (c, d) shows the true energy landscape (i.e., $-\log p_{\mathcal{M}}$) on the UCG and WCG datasets, respectively.
@@ D.3 Quantitative evaluation with error bars
Metric $p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$ ()($\uparrow$)( ↑ ) RMSE ()($\downarrow$)( ↓ )
$\boldsymbol{G}_{E_{\mathcal{M}}}$ $0.79\pm 0.02$
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ $0.77\pm 0.04$
[…]
@@
[…]
@@
Metric $p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$ ()($\uparrow$)( ↑ ) RMSE ()($\downarrow$)( ↓ )
$\boldsymbol{G}_{E_{\mathcal{M}}}$ $0.67\pm 0.05$
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ $0.73\pm 0.07$
[…]
@@
[…]
@@ E.1 Datasets
In this setting, character orientations are sampled uniformly across the full range of $[-179^{\circ},180^{\circ}]$, using a one-degree step. This ensures that each possible orientation within this interval is equally likely. Importantly, the distribution is consistent across all characters, meaning that each character appears with the same uniform spread of rotations.
@@ E.2 Architecture and algorithm of the Triplet Loss autoencoder
       Sample bath of triplet B=$(\mathbf{x}_{a},\mathbf{x}_{p},\mathbf{x}_{n})$ from $\mathcal{D}$
@@ E.5 Additional geodesics
Refer to captiona)${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$b)${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$c)${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$d)${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$e) Linear interpolation
@@
Refer to captiona)${\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}}$b)${\color[rgb]{0.91796875,0.26171875,0.20703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.91796875,0.26171875,0.20703125}\mathbf{G_{RBF}}}$c)${\color[rgb]{0.203125,0.66015625,0.32421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.66015625,0.32421875}\mathbf{G_{1/p_{\theta}}}}$d)${\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}}$e) Linear interpolation
@@ E.6 Quantitative evaluation with error bars
Metric $\mathcal{D}$-RMSE ()($\downarrow$)( ↓ ) $\gamma^{\star}$-RMSE ()($\downarrow$)( ↓ )
linear interp. $2.96\pm 0.42$ $3.52\pm 0.21$
[…]
@@
[…]
@@ 1 Conv2d ($8F$, 1, 2, 1, 0): for $f_{3}$
Table 9: Architecture of the energy function. $F$ denotes the base number of feature channels, and $n_{c}$ is the number of input channels. The final energy is computed using three parallel output heads. The notation Conv2d($n_{c}$, $n_{f}$, $k$, $s$, $p$) refers to a 2D convolutional layer with $n_{c}$ input channels, $n_{f}$ output channels, a kernel size of $k$, stride $s$, and padding $p$.
@@ F.4 FIDs with error bars
[…]
@@
[…]
@@ F.6 About the Spherical interpolation
Figure 25: Distribution of $||\mathbf{x}||^{2}$ on the AFHQ dataset
This interpolation method, introduced by [60], has proven particularly effective for interpolating in the latent space of VAEs. The intuition behind its success is that it implicitly assumes the data manifold lies on a hypersphere. While this may seem restrictive, the assumption is reasonable in practice. In a VAE, each latent coordinate $x_{i}$ is drawn from a standard Normal distribution: $x_{i}\sim\mathcal{N}(0,1)$ ($1<i<D$). As a result, the squared norm, $||\mathbf{x}||^{2}=\sum_{i=1}^{D}x_{i}^{2}$ follows a chi-squared distribution with $D$ degree of freedom. This distribution is known to concentrate tightly around $D$, effectively placing most latent codes near the surface of a hypersphere. To validate this empirically, we visualize the distribution of $||\mathbf{x}||^{2}$ for all latent codes of the AFHQ dataset (see Fig. 25). We observe that this distribution is concentrated on $D$$D=1024$ for VAE with latent space of size $4\times 16\times 16$.
@@ F.7 Physical interpretation
[…]
@@
  • $-\,\frac{1}{2}\|\dot{\boldsymbol{\gamma}}(t)\|^{2}\cdot\nabla\log p(\boldsymbol{\gamma}(t))$ is a force in the direction of the stein score (pointing toward low density regions). It behaves like a repulsive force, pushing the particle toward areas with low probability. The faster the particle moves, the stronger the force.
@@
This equation can be interpreted through Newton’s second law: it describes the motion of a particle $\mathbf{x}$ following a geodesic in the Riemanannian manifold $\big{(}\mathcal{M},\frac{1}{p(\mathbf{x})}\big{)}$, where $p(\mathbf{x})$ denotes the data density. The particle experiences a force $-\frac{1}{2}\|\dot{\mathbf{x}}\|^{2}\nabla\log p(\mathbf{x})$, pushing away from regions of high probability. The term $||\mathbf{x}||^{2}$ modulates the forces magnitude and plays a role analogous to momentum, strengthening the pull when the particle moves quickly. While this is not a literal physical system—here the particle is a data point, and has no mass, it provides a useful analogy for understanding the dynamics of trajectories shaped by data geometry.

### v3

@@ 4.3 Rotated Characters
At first glance, it may seem counterintuitive that trajectories following the geodesics move faster in high-density regions. However, this is consistent with minimizing the kinetic energy $\mathcal{E}$ in Eq.2, which enforces constant Riemannian speed (i.e., the quantity $\dot{\mathbf{\gamma}}(t)||_{\mathbf{\gamma}(t)}$ is preserved along the trajectory) but not a constant Euclidean speed (i.e., $\dot{\mathbf{\gamma}}(t)||$ is not constant). Since EBM-derived metrics assign lower Riemannian cost in high-density regions, maintaining constant Riemannian speed requires moving faster in Euclidean terms through these regions. The faster rotation near 0°, observed in Fig.5 and Fig. 3, thus reflects the lower Riemannian cost of traveling through high-density regions. These results confirm and extend our previous findings: metrics based on energy (i.e., proportional to $\log p$) more effectively capture the curvature of the data manifold.
@@ 4.4 Animal Faces
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} 20.79
@@
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} 39.17
@@
Table 1: FID along geodesics for different Riemannian metrics. FID is computed at each trajectory point to assess on-manifold alignment. See Supp. F.4 for the 22-$\sigma$ error.
@@ 5 Conclusion
We focus on conformal metrics, which scale the identity by a scalar field to encode density. While more complex, non-conformal and anisotropic metrics (e.g., the Stein metric [25]) are accessible from the EBM score, we found that conformal metrics offer comparable performance with simpler interpretation and reduced computational cost, justifying our focus in this work. Future work may explore these extensions with regularization or structural priors to ensure smoothness and scalability (See Supp.G for a discussion of limitations and Supp.H for broader impact). To keep computational cost manageable, we train the EBM in the latent space of a pretrained autoencoder and compute geodesics using finite-difference optimization, two design choices that substantially reduce complexity and memory use without compromising performance.
Although this article is primarily methodological, it points to promising applications. One example is the mental rotation task, in which humans mentally rotate objects to match a target [62]. In such experiments, reaction times tend to decrease with training [63], suggesting that repeated exposure sharpens internal representations around training examples. These refined representations may concentrate in high-density regions, where mental transformations occur more quickly. As shown in Fig. 3 and 5, our geodesics naturally accelerate in such high-density regions, echoing these psychophysical findings. Modeling mental simulation as geodesics on Riemannian manifolds shaped by a generative model offers a principled computational framework to understand human cognition. It provides a way to formalize and test the hypothesis that the human cognition relies on generative models to support flexible inference [64, 65, 66, 67, 68]. Our approach is also particularly relevant for neuroscience, where datasets are high-dimensional, often sparsely sampled, and where understanding the geometry of neural population activity is central to scientific insight. In such settings, high-fidelity geodesics are essential for capturing the true structure of neural trajectories—approximations may distort the manifold and lead to misinterpretation of brain dynamics. While training EBMs is costly, the benefits in terms of interpretability and geometric accuracy make this approach compelling for applications where precision is critical.
@@
Acknowledgments
This work was supported by ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A-0004). Part of this work was carried out within the DEEL project, which is part of IRT Saint Exupéry and the ANITI AI cluster. The authors acknowledge the financial support from DEEL’s industrial and academic members and the “France 2030” program (NR-10-AIRT-01 and ANR-23-IACL-0002). Additional support for TS provided by ONR (N00014-24-1-2026 and REPRISM MURI N00014-24-1-2603) and NSF (IIS-2402875).
@@ B.1 Derivation of the Gradient of the EBM Log-Likelihood
$p_{\theta}(\mathbf{x})=\frac{\exp(-E_{\theta}(\mathbf{x}))}{Z(\theta)}\quad\text{with}\quad Z(\theta)=\int\exp(-E_{\theta}(\mathbf{x}))\,d\mathbf{x}.$
@@
$\log p_{\theta}(\mathbf{x})=E_{\theta}(\mathbf{x})+\log Z(\theta).$
@@
$=\frac{1}{Z(\theta)}\nabla_{\theta}\int\exp(-E_{\theta}(\mathbf{x}))\,d\mathbf{x}$
$=-\frac{1}{Z(\theta)}\int\exp(-E_{\theta}(\mathbf{x}))\nabla_{\theta}E_{\theta}(\mathbf{x})\,d\mathbf{x}$
$=-\int p_{\theta}(\mathbf{x})\nabla_{\theta}E_{\theta}(\mathbf{x})\,d\mathbf{x}$
@@ B.2 EBM training algorithm
Input: Training dataset $\mathcal{D}$, learning rate $\eta$, Replay Buffer $\mathcal{B}$, Langevin step size $\alpha$, noise scale $\sigma$, number of Langevin steps $L$
@@
 $\mathbf{x}^{-}=\mathbf{x}^{L}\textnormal{.detach()}$
 $\nabla_{\theta}\mathcal{L}_{\text{ML}}\approx\mathbb{E}_{\textnormal{Batch}}\left[\nabla_{\theta}E_{\theta}(\mathbf{x}^{+}_{i})-\nabla_{\theta}E_{\theta}(\mathbf{x}^{-}_{i})\right]$ ## Compute the ML loss
 $\mathcal{L}_{REG}(\theta)=\mathbb{E}_{\textnormal{Batch}}\left[\nabla_{\theta}E_{\theta}(\mathbf{x}^{+}_{i})^{2}+\nabla_{\theta}E_{\theta}(\mathbf{x}^{-}_{i})^{2}\right]$ ## Compute Regularization loss
 $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{ML}}-\eta\nabla_{\theta}\mathcal{L}_{REG}$ ## update parameters with gradient descent
@@ C.2 LAND metric
[…]
@@ C.3 RBF metric
[…]
@@ D.2 Neural networks architectures and Hyperparameters on the Circular Mixture of Gaussian Dataset
[…]
Figure 10: Energy Landscape on the UCG and WCG datasets. (a, b) shows the energy landscape learned by the EBMs on the UCG and WCG datasets, respectively. (c, d) shows the true energy landscape (i.e., $\log p_{\mathcal{M}}$) on the UCG and WCG datasets, respectively.
@@ D.3 Quantitative evaluation with error bars
Metric
$p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$
$(\uparrow)$
RMSE
$(\downarrow)$
$\boldsymbol{G}_{E_{\mathcal{M}}}$ $0.79\pm 0.02$ -
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ $0.77\pm 0.04$ -
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} $\mathbf{0.78}\pm 0.03$ $0.12\pm 0.02$
@@
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} $0.60\pm 0.07$ $0.38\pm 0.05$
@@
Metric
$p_{\mathcal{M}}(\mathbf{\gamma}^{\star})$
$(\uparrow)$
RMSE
$(\downarrow)$
$\boldsymbol{G}_{E_{\mathcal{M}}}$ $0.67\pm 0.05$ -
$\boldsymbol{G}_{1/p_{\mathcal{M}}}$ $0.73\pm 0.07$ -
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} $\mathbf{0.67}\pm 0.06$ $0.18\pm 0.07$
@@
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} $0.65\pm 0.11$ $0.34\pm 0.05$
@@ E.1 Datasets
In this setting, character orientations are sampled uniformly across the full range of $179^{\circ},180^{\circ}]$, using a one-degree step. This ensures that each possible orientation within this interval is equally likely. Importantly, the distribution is consistent across all characters, meaning that each character appears with the same uniform spread of rotations.
@@ E.2 Architecture and algorithm of the Triplet Loss autoencoder
[…]
@@ E.5 Additional geodesics
[…]
@@
[…]
@@ E.6 Quantitative evaluation with error bars
Metric
$\mathcal{D}$-RMSE
$(\downarrow)$
$\gamma^{\star}$-RMSE
$(\downarrow)$
linear
interp.
$2.96\pm 0.42$ $3.52\pm 0.21$
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} $\mathbf{0.11}\pm 0.01$ $\mathbf{0.40}\pm 0.03$
@@
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} $0.66\pm 0.12$ $2.39\pm 0.51$
@@ 1 Conv2d ($8F$, 1, 2, 1, 0): for $f_{3}$
Table 9: Architecture of the energy function. $F$ denotes the base number of feature channels, and $n_{c}$ is the number of input channels. The final energy is computed using three parallel output heads. The notation Conv2d($n_{c}$, $n_{f}$, $k$, $s$, $p$) refers to a 2D convolutional layer with $n_{c}$ input channels, $n_{f}$ output channels, a kernel size of $k$, stride $s$, and padding pp.
@@ F.4 FIDs with error bars
$\color[rgb]{0.2578125,0.5234375,0.95703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2578125,0.5234375,0.95703125}\mathbf{G_{E_{\theta}}}$} $20.79\pm 2.17$
@@
$\color[rgb]{0.984375,0.73828125,0.01953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.984375,0.73828125,0.01953125}\mathbf{G_{LAND}}$} $39.17\pm 3.63$
@@ F.6 About the Spherical interpolation
Figure 25: Distribution of $\mathbf{x}||^{2}$ on the AFHQ dataset
[…]
@@ F.7 Physical interpretation
where $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ are the usual Euclidean inner product and norms, respectively.
@@
  • $\frac{1}{2}\|\dot{\boldsymbol{\gamma}}(t)\|^{2}\cdot\nabla\log p(\boldsymbol{\gamma}(t))$ is a force in the direction of the stein score (pointing toward low density regions). It behaves like a repulsive force, pushing the particle toward areas with low probability. The faster the particle moves, the stronger the force.
@@
This equation can be interpreted through Newton’s second law: it describes the motion of a particle $\mathbf{x}$ following a geodesic in the Riemanannian manifold $\big(\mathcal{M},\frac{1}{p(\mathbf{x})}\big)$, where $p(\mathbf{x})$ denotes the data density. The particle experiences a force $\frac{1}{2}\|\dot{\mathbf{x}}\|^{2}\nabla\log p(\mathbf{x})$, pushing away from regions of high probability. The term $\mathbf{x}||^{2}$ modulates the forces magnitude and plays a role analogous to momentum, strengthening the pull when the particle moves quickly. While this is not a literal physical system—here the particle is a data point, and has no mass, it provides a useful analogy for understanding the dynamics of trajectories shaped by data geometry.